# A Comparison of Relativistic Coupled Cluster and Equation of Motion Coupled Cluster Quadratic Response Theory

### v1

@@ 2.2 Implementation details
Our QR-CC implementation has been carried out in the ExaCorr module 84 of the relativistic quantum chemistry program DIRAC127. Below we provide details on the implementation of the QR-CC response function calculation, and refer the reader to prior works concerning (a) the solution of ground-state amplitudes and multipliers 84; and (b) the solution of first-order response equations 113, which leveraged our prior implementation of a generalized Davidson eigenvalue solver128, 102 to implement the solution of systems of linear equations 129. We recall that through the use of the TAL-SH tensor library 130, a modular component of the ExaTENSOR library 131, all tensor operations in our code are parallelized via OpenMP and can be offloaded to NVIDIA or AMD GPUs whenever these are available.
@@
In our implementation, we divide the calculation into different levels of subroutines: (a) a top-level driver to organize calculations for and and to apply the symmetrization operator as in Eq. 12; (b) a driver that calculates for all operators under consideration, solves all corresponding linear response equations for either QR-CC and QR-EOMCC, and calls the driver to calculate the appropriate response function.
@@
In the case of QR-CC, the response function driver carries out the six permutations between the triplet of operators; it calls a subroutine that calculates the terms within the curly braces in Eq.13 for a given pair or operators and stores it on set of 2-/4-dimension tensors (singles and doubles manifolds) and performs the contraction with the perturbed amplitudes. This allows us to, if desired, extract the contributions from each of the four terms (for each triplet of operators) and further break these down in terms of the contributions from the singles and doubles manifolds.
@@ 3 Computational details
All QR-CC, QR-EOMCC calculations were carried out with a development version (revision number 8ef365b7) of the DIRAC code127, 132. All coupled cluster calculations have been carried out in GPU-accelerated nodes.
@@
As done in our prior EOM-QR calculations 115, we have used some of the functionality connected to the Cholesky-decomposition approach in ExaCorr84 to avoid the explicit construction of the 2-electron repulsion integral tensor in AO basis in the integral transformation step and with that reduce the memory footprint of our calculations. We employed a thresholds of 10-5 for all systems except HAt and HTs, for which the threshold was set to 10-4. The outcome of response calculations is not significantly affected by these approximations.
@@
The Dunning basis set 133, 134, 135 augmented by three diffuse functions (denoted by t-aug-cc-pVTZ) was used for the hydrogen atom. For all other atoms, the Dyall triple zeta sets (denoted by t-aug-dyall.v3z) were employed in general, though in some cases and for elements up to bromide the t-aug-cc-pVTZ set was used to make a direct comparison to our prior EOM-QR calculations 115 possible. The diffuse functions in question were automatically generated by DIRAC as an even-tempered sequence based on the most diffuse exponents for all angular momenta primitives in the original basis.
@@
In addition to the non-relativistic 136, 137(as activated by the .Levy-Leblond keyword) and eXact 2-Component (X2C) relativistic123, in the coupled cluster calculations we also investigated the molecular mean-field X2C method based on the Dirac-Coulomb (2DCM) and Dirac-Coulomb-Gaunt (2DCGM) Hamiltonians, as well as the spin-free (SF) variants of the X2C and 2DCGM Hamiltonians. In what follows, we shall use the term orbital as shorthand for both spinors and spin-orbitals, depending on the Hamiltonian used in the calculation. In X2C calculations including spin-orbit coupling, we make use of atomic-mean field SOC integrals, obtained from a calculation employing the Douglas-Kroll-Hess Hamiltonian.
@@
The relativistic and non-relativistic calculations have been carried out with the Gaussian type138 and point charge nucleus model, respectively.
@@
The molecular structures employed in our calculations have been taken from Huber and Herzberg 139 for HX (X=F, Cl, Br, I), from Gomes and Visscher 140 for HAt, and from de Farias 141 for HTs, so respectively: H-F (0.91680 Å), H-Cl (1.27455 Å), H-Br (1.41443 Å), H-I (1.60916 Å), H–At (1.722 Å) and H-Ts (1.9676 Å). As the choice of orientation in these systems may affect the signs of the electric dipole and first hyperpolarizability (but not their absolute value), we note our results have been tabulated following the convention that the heavy atoms are placed at the origin of the coordinate system, with the molecular axis aligned with the z axis of the coordinate system.
For the most part we retained the correlation spaces defined by Yuan et al. 115, which include valence electrons and virtual orbitals with energies up to 5 a.u.; However, in some cases (Xe, Rn, Og) we also explored the correlation of the d shell and in selected cases, larger virtual cutoffs (100 a.u.), and in others (HF, HCl, HBr) we also carried out calculations correlating all electrons and all virtual orbitals.
@@ 4 Results and discussion
In the sections below we discuss our results for the first hyperpolarizability and Verdet constant for systems across the periodic table. We note that the choice of computational setup concerning the basis sets and correlation space, as well as a comparison to the B3LYP functional, have been discussed in detail in our work discussing the implementation of QR-EOMCC 115. Therefore, in the following we will mostly focus on a comparison between QR-CC and QR-EOMCC for different Hamiltonians.
@@ 4.1 First hyperpolarizability of hydrogen halides (HX, X=F–Ts)
Due to the symmetry of the systems under consideration, we restrict ourselves to the calculation of the parallel component of the static first hyperpolarizability ()142
where the element is given by the sum-over-states expression1:
with the Cartesian components of the electric dipole operators, indicating the sum of six terms by permuting the pairs (), (), (), and denoting the excitation energy between the ground state and excited state . In the following we only calculated two unique components, and , so that and are calculated from
respectively. Our results for the frequency-independent case are summarized in tables 2 and 3 for the and . To illustrate the trends for the individual components, in figure 1 we present results for the SF-2DCM and 2DCM Hamiltonians, while in tables LABEL:SI-tab:hyperpolar-table1 and LABEL:SI-tab:hyperpolar-table2 in the SI we report the values for the and components respectively for all Hamiltonians under consideration.
We recall that our use of a Gaussian nuclear model and Dyall bases in all calculations makes for a slightly different setup than that of Yuan et al. 115. For ease of comparison, we also provide results obtained with the Dunning basis set on the halogens, and note small discrepancies for calculations with the Levy-Leblond Hamitonian arising from the difference in nuclear model.
@@
We have noticed small differences between our results and those previously reported 115, the latter showing larger (in absolute value) hyperpolarizabilties. We have traced this difference to the use of Dyall basis sets for the current calculations, as opposed to the Dunning basis sets. This difference can be understood as a consequence of the bias towards somewhat less diffuse exponents as part of the process of obtaining relativistic energy-optimized basis sets.
@@
Quantifying the differences between QR-CC and QR-EOMCC by computing (), we see that differences are small for HF but become non-negligible already for HCl and remain sizable for the other species. In relative terms, the differences for are about 6% for HF, but grow to about 30% for HCl, 120 % for HBr and 150% for HI. For the very large positive hyperpolarizability of HTs, the difference is in absolute sense the largest, but amounts to only 15% of the 2DCGM value. For the more accessible HAt molecule, the effect is similar but less pronounced. Finally, we see essentially the same trends when correlating more electrons and virtuals in the lighter systems (HF to HBr), though we do observe a somewhat smaller difference between QR-CC and QR-EOMCC for HBr than when only the valence is correlated. This is driven by a larger increase in the for QR-EOMCC than for QR-CC.
We note the situation is a bit different for , where we observe that apart from HTs, the difference between is QR-CC and QR-EOMCC is fairly modest across the series.
@@
Going beyond the static case, in figure 2 we present the results for second harmonic generation (SHG) calculations on HI, which corresponds to the case in which . As this property has been discussed in detail in our work presenting QR-EOMCC 115, we refer the reader to it for further details on the behavior of the dispersion curves. We nevertheless highlight the fact that that in the region around a.u., corresponding to half the excitation energy of the spin–orbit allowed transition towards the state, we see the expected pole for the quadratic response functions. Beside that, we observe that across the range of frequencies sampled, the difference between QR-CC and QR-EOMCC results amounts to a constant shift.
@@
In the following we look more closely at the differences between QR-CC and QR-EOMCC by analyzing the contributions from the different terms to the response functions, found in tables LABEL:SI-tab:breakdown-qrf-eomcc and LABEL:SI-tab:breakdown-qrf-cc for QR-CC and QR-EOMCC respectively, as we go across the series from HF to HI.
@@
For this analysis, it is interesting to also consider the results for the static dipole polarizabilties for the same molecules, since the difference between LR-CC and LR-EOMCC calculations is connected to the presence of the term in the LR-CC response function (and therefore the right-hand side transition moment), whereas LR-EOMCC employs a simpler expression (for both methods perturbed amplitudes are obtained from eq. 10a).
@@
These dipole polarizabilities, shown in figure 3 (values for dipole moments and polarizabilities are reported in the SI) increase across the series, which is consistent with the increase in size of the halides accompanied by their decreasing electronegativity. The differences between CC and EOMCC do increase across the series for the and components and consequently, for the mean polarizability as well. However, in contrast to the hyperpolarizabilities, in relative terms these differences are roughly the same and of about 2% for and 1.5% for .
@@
The difference between CC and EOMCC is more strongly felt for the anisotropy, which is about 3% for HF, but 8-12% for HCl to HAt, and reaches about 30% for HTs, though in absolute terms the differences are still modest. We note that our CCSD dipole moments compare well with available experimental 143, 144 or theoretical work (for HTs) 145, though for reaching benchmark accuracy we would need to account for higher excitations, something which is beyond the scope of this work. These findings add to the body of data supporting the argument that differences between LR-CC and LR-EOMCC properties are in practice small.
@@
Considering now the individual contributors to the value of the QR-CC functions, we see that dominant contributions typically arise from the terms from eq. 14b, involving both the perturbed amplitudes and multipliers ( and ). Furthermore, we observe that these two contributions to the final value in general tend to counteract each other, and that the magnitude of the “doubles” contributions increases as the halide becomes heavier. Terms from eq. 14a, involving only perturbed amplitudes and ground-state multipliers ( and ) show smaller (but non-negligible) net contributions than those above. They also increase in absolute value as the halide becomes heavier.
@@
The relative importance of contribution of the terms in eq. 14b also changes depending on which response function we look at. For , the term is the one which contributes the most, but for there are more important cancellations between the “singles” and “doubles” contributions from it so that the term becomes a more important net contributor, being the dominant one for the HI molecule followed by , whereas and show similar contributions.
@@
4.2 Verdet constant of Xe, Rn and Og

### v3

@@ 2.2 Implementation details
Our QR-CC implementation has been carried out in the ExaCorr module pototschnig_implementation_2021 of the relativistic quantum chemistry program DIRACsaue2020dirac. Below we provide details on the implementation of the QR-CC response function calculation, and refer the reader to prior works concerning (a) the solution of ground-state amplitudes and multipliers pototschnig_implementation_2021; and (b) the solution of first-order response equations Yuan_Xiang_LR, which leveraged our prior implementation of a generalized Davidson eigenvalue solverhirao1982generalization, shee2018equation to implement the solution of systems of linear equations olsen1988solution. We recall that through the use of the TAL-SH tensor library githubRepo2023, a modular component of the ExaTENSOR library lyakh_domainspecific_2019, all tensor operations in our code are parallelized via OpenMP and can be offloaded to NVIDIA or AMD GPUs whenever these are available.
@@
In our implementation, we divide the calculation into different levels of subroutines: (a) a top-level driver to organize calculations for and and to apply the symmetrization operator as in Equation 10k; (b) a driver that calculates for all operators under consideration, solves all corresponding linear response equations for either QR-CC and QR-EOMCC, and calls the driver to calculate the appropriate response function.
@@
In the case of QR-CC, the response function driver carries out the six permutations between the triplet of operators; it calls a subroutine that calculates the terms within the curly braces in subsection 2.1 for a given pair or operators and stores it on set of 2-/4-dimension tensors (singles and doubles manifolds) and performs the contraction with the perturbed amplitudes. This allows us to, if desired, extract the contributions from each of the four terms (for each triplet of operators) and further break these down in terms of the contributions from the singles and doubles manifolds.
The correctness of our QR-CC implementation was validated through a comparison of non-relativistic results (for the HF, HCl and HBr systems) with (a) the analytic implementation of QR-CC in the Dalton aidas2014d, Dalton2020p1 code (for , see LABEL:SI-tab:verification-analytic-hf, LABEL:SI-tab:verification-analytic-hcl and LABEL:SI-tab:verification-analytic-hbr); and (b) finite-field calculations with the ExaCorr code, where the z-component of the electric dipole was employed as perturbing operator (with different perturbation strenghts) and included to the coupled cluster energy calculations. With that, we ensure that the numerical derivatives values (for , see LABEL:SI-tab:compare-nrdz-hf-hbr) are consistent with the orbital-unrelaxed formulation of CC response. We also employed the analytic implementation of LR-CC in the CFOUR code matthews2020coupled to provide an additional check for , and to investigate the relative importance of orbital relaxation for these systems (at the time of writing the QR-CC functionality in CFOUR is not part of the public release). In these calculations, carried out for HF, HCl and HBr, we employed contracted Dunning basis sets, a point-charge nucleus, and correlated all electrons and virtual orbitals. We note that we have at the same time validated the correctness of the code to calculate the left and right-hand side quadratic response equations, by calculating the components of the dipole polarizability with the asymmetric formulation of linear response theory (see e.g. Yuan_Xiang_LR) using the perturbed amplitudes and multipliers that go in the calculation of the QR-CC response functions. The results and computational procedures associated to these validations are found in the SI.
Our code currently lacks the functionality to obtain QR-EOMCC molecular properties from finite-field calculations with the EOMCC model, and implementing such functionality is beyond the scope of this work. As explained in Yuan_Xiang_QR, we had verified our QR-EOMCC implementation against static values reported by coriani_molecular_2016, obtained with a QR-EOMCC implementation in the Dalton code (which, at the time of writing, is not publicly available). For convenience, a comparison between codes for the HF molecule is presented in the SI.
@@ 3 Computational details
All QR-CC, QR-EOMCC calculations were carried out with a development version (revision number 8ef365b7) of the DIRAC codesaue2020dirac, DIRAC25. All coupled cluster calculations have been carried out in GPU-accelerated nodes.
@@
As done in our prior EOM-QR calculations Yuan_Xiang_QR, we have used some of the functionality connected to the Cholesky-decomposition approach in ExaCorrpototschnig_implementation_2021 to avoid the explicit construction of the 2-electron repulsion integral tensor in AO basis in the integral transformation step and with that reduce the memory footprint of our calculations. We employed a threshold of 10-5 for all systems except HAt and HTs, for which the threshold was set to 10-4. The outcome of response calculations is not significantly affected by these approximations.
@@
In ref. Yuan_Xiang_QR, which describes our QR-EOMCC implementation, we investigated for the HF molecule the effect of basis set quality and virtual space truncation on the hyperpolarizabilities, and arrived at the conclusion that with a setup consisting of triply augmented triple-zeta basis sets (the Dunning basis set kendall1992electron, woon1993gaussian, wilson1999gaussian for H and the Dyall basis sets dyall2006relativistic, basis-Dyall-TCA2012-131-3962 for all other atoms) was sufficiently converged for investigating the trends across the periodic table. In some cases and for elements up to bromide the t-aug-cc-pVTZ set was used to make a direct comparison to our prior EOM-QR calculations Yuan_Xiang_QR possible. The diffuse functions in question were automatically generated by DIRAC as an even-tempered sequence based on the most diffuse exponents for all angular momenta primitives in the original basis. In this manuscript we provide additional elements to show this choice of basis set was justified by carrying out a basis set study for the HF, HCl and HBr molecules.
@@
In addition to the non-relativistic Levy-Leblond Hamiltonian (LL) levy1967nonrelativistic, visscher2000approximate(as activated by the .Levy-Leblond keyword) and eXact 2-Component (X2C) relativisticiliavs2007infinite, in the coupled cluster calculations we also investigated the molecular mean-field X2C method based on the Dirac-Coulomb (2DCM) and Dirac-Coulomb-Gaunt (2DCGM) Hamiltonians, as well as the spin-free (SF) variants of the X2C and 2DCGM Hamiltonians. In what follows, we shall use the term orbital as shorthand for both spinors and spin-orbitals, depending on the Hamiltonian used in the calculation. In X2C calculations including spin-orbit coupling, we make use of atomic-mean field SOC integrals, obtained from a calculation employing the Douglas-Kroll-Hess Hamiltonian.
For the correlation spaces, these included the valence electrons (for a total 8 electrons corresponding to those from the 1s shell of hydrogen, the ns2np5 shells for the halides and the ns2np6 shells for noble gases) and virtual orbitals with energies up to 5 a.u., a value that was also established in ref. Yuan_Xiang_QR on the basis of calculations for HF. To illustrate the effect of this choice on the results, For HF, HCl and HBr we have also explored larger virtual cutoffs (10 and 100 a.u.) and no cutoffs (all virtuals correlated), and carried out calculations correlating all electrons and all virtual orbitals employing the 2DCM Hamiltonian. Additionally, in the case of the noble gases (Xe, Rn, Og) we also explored the correlation of the d shell.
@@
The relativistic and non-relativistic calculations have been carried out with the Gaussian typevisscher1997dirac and point charge nucleus model, respectively.
@@
The molecular structures employed in our calculations have been taken from huber1979constants for HX (X=F, Cl, Br, I), from pereira_gomes_influence_2004 for HAt, and from deFarias2017 for HTs, so respectively: H-F (0.91680 Å), H-Cl (1.27455 Å), H-Br (1.41443 Å), H-I (1.60916 Å), H–At (1.722 Å) and H-Ts (1.9676 Å). As the choice of orientation in these systems may affect the signs of the electric dipole and first hyperpolarizability (but not their absolute value), we note our results have been tabulated following the convention that the heavy atoms are placed at the origin of the coordinate system, with the hydrogen atom placed on the positive z-axis of the coordinate system.
@@ 4 Results and discussion
In the sections below we discuss our results for the first hyperpolarizability and Verdet constant for systems across the periodic table. We note that the choice of computational setup concerning the basis sets and correlation space, as well as a comparison to the B3LYP functional, have been discussed in detail in our work discussing the implementation of QR-EOMCC Yuan_Xiang_QR. Therefore, in the following we will mostly focus on a comparison between QR-CC and QR-EOMCC for different Hamiltonians.
@@ 4.1 First hyperpolarizability of hydrogen halides (HX, X=F–Ts)
Due to the symmetry of the systems under consideration, we restrict ourselves to the calculation of the parallel component of the static first hyperpolarizability ()shelton_measurements_1994
where the element is given by .
In the following we only calculated two unique components, and , so that and are calculated from
respectively. Our results for the frequency-independent case are summarized in Table 2 for and in Table 3 for . To illustrate the trends for the individual components, in Figure 1 we present results for the SF-2DCM and 2DCM Hamiltonians, while in LABEL:SI-tab:hyperpolar-table1 and LABEL:SI-tab:hyperpolar-table2 we report the values for the and components respectively for all Hamiltonians under consideration.
We recall that our use of a Gaussian nuclear model and Dyall bases in all calculations makes for a slightly different setup than that of Yuan_Xiang_QR. For ease of comparison, we also provide results obtained with the Dunning basis set on the halogens, and note small discrepancies for calculations with the Levy-Leblond Hamitonian arising from the difference in nuclear model.
@@
We have noticed small differences between our results and those previously reported Yuan_Xiang_QR, the latter showing larger (in absolute value) hyperpolarizabilties. We have traced this difference to the use of Dyall basis sets for the current calculations, as opposed to the Dunning basis sets. This difference can be understood as a consequence of the bias towards somewhat less diffuse exponents as part of the process of obtaining relativistic energy-optimized basis sets.
@@
Quantifying the differences between QR-CC and QR-EOMCC by computing (), we see that differences are small for HF but become non-negligible already for HCl and remain sizable for the other species. In relative terms, the differences for are about 6% for HF, but grow to about 30% for HCl, 120 % for HBr and 150% for HI. For the very large positive hyperpolarizability of HTs, the difference is in absolute sense the largest, but amounts to only 15% of the 2DCGM value. For the more accessible HAt molecule, the effect is similar but less pronounced. We note the situation is a bit different for , where we observe that apart from HTs, the difference between is QR-CC and QR-EOMCC is fairly modest across the series.
The results above have been obtained with a computational setup (triply augmented basis sets and virtual spaces truncated at 5 a.u.) established on the basis of a verification of basis set and correlation effects for the HF molecule in our paper describing the implementation of QR-EOMCC Yuan_Xiang_QR, and aimed at consistently addressing the whole hydrogen halide series. It is nevertheless important to have an idea of how close this setup is to the converged basis set and correlation treatment. To this end, we have investigated the effect of varying the number of diffuse functions and the cardinal number of the basis sets for HF, HCl and HBr, employing the 2DCM Hamiltonian.
From our results, found in LABEL:SI-tab:InfluenceOfDiffuseSI, we have that for triple-zeta basis sets, the triply and quadruply augmented basis sets typically show agreement better than 0.1 a.u. for all molecules, effectively validating our choice and in line with prior work on the light hydrogen halides Fernandez1998. We consider nevertheless that doubly augmented triple-zeta basis set can still be a good choice whenever it is not possible to employ larger basis sets for heavier elements. Furthermore, as expected singly augmented triple zeta bases are not sufficiently flexible, but are definitely much better suited than any double-zeta basis.
Another surprising but very interesting behavior revealed by this benchmark is that QR-CC and QR-EOMCC results are actually very close to each other for double-zeta basis sets without diffuse functions–including for HCl and HBr–and as soon as diffuse functions are added the differences between the two methods start to grow. To the best of our knowledge, this point has not been explored in the literature in detail, and possibly is the reason why our results seem at first at odds with the claims in the literature that CC and EOMCC yield results which are typically very close to each other.
With respect to the convergence of the correlation space, we observe from Table 2 and Table 3 that the 5 a.u. threshold is not a poor choice, but it is clear that with it the virtual space is not converged. Upon exploring other truncations while correlating only the valence electrons, we see that at 10 a.u. results are not quite converged either, whereas with a truncation at 100 a.u. results are essentially equivalent to those correlating all virtuals. An important finding is that, in qualitative terms, there is no difference in the trends across the halogen series for all virtual truncations employed. Finally, for HBr core correlation is not negiglible, accounting for an increase of 4% and 18% for calculated with QR-CC and QR-EOMCC. This change reduces the difference between the two approaches by nearly 1 a.u., but in any case the discrepancy remains significant (larger than 4 a.u.). It is important to keep in mind that such change is driven primarily by an increase in .
@@
Going beyond the static case, in Figure 2 we present first the results for second harmonic generation (SHG) calculations on HI, which corresponds to the case in which . As this property has been discussed in detail in our work presenting QR-EOMCC Yuan_Xiang_QR, we refer the reader to it for further details on the behavior of the dispersion curves. We nevertheless highlight the fact that that in the region around a.u., corresponding to half the excitation energy of the spin–orbit allowed transition towards the state, we see the expected pole for the quadratic response functions. Beside that, we observe that across the range of frequencies sampled, the difference between QR-CC and QR-EOMCC results amounts to a nearly constant shift for low frequencies. For the region close to the resonance, on the other hand, we see a relatively small decrease in the difference between QR-CC and QR-EOMCC.
In Table 4 we present a second dynamic property, the electro-optical Pockels effect boyd2020nonlinear, norman_principles_2018, which corresponds to the condition . The Pockels effect can be understood as capturing the change in the dynamic polarizability due to an applied static electric field. A first consequence of the condition is that the resonances now only occur near the first excited states (e.g. for HI, at about a.u.) instead of at smaller frequencies such as in SHG. This means that for the systems under study we are very much off-resonance and consequently the values remain very similar to those in the static case. Interestingly, we see a (very small) decrease in the difference between QR-CC and QR-EOMCC as the perturbation frequency increases, as found for the SHG case, but that in no way is capable of reversing the growing difference between QR-CC and QR-EOMCC across the halide series, and that these differences are driven by the component.
@@
In the following we look more closely at the differences between QR-CC and QR-EOMCC by analyzing the contributions from the different terms to the response functions, found in SI (LABEL:SI-tab:breakdown-qrf-cc-beta for QR-CC and LABEL:SI-tab:breakdown-qrf-eomcc-beta for QR-EOMCC respectively), as we go across the series from HF to HI.
@@
For this analysis, it is interesting to also consider the results for the static dipole polarizabilties for the same molecules. The reason for that is twofold: first, it is well-established that the formal difference between LR-CC and LR-EOMCC is connected to the presence of the term in the LR-CC response function (and therefore the right-hand side ground to excited state transition matrix elements), whereas LR-EOMCC employs a simpler expression–for both methods perturbed amplitudes are obtained from eq. Equation 10a, and left-hand side transition matrix elements are formally the same. This means that a large difference in polarizability is indicative of the relative importance of the full term, and consequently of an important difference in the right-hand side ground-to-excited state transition matrix elements.
Second, if one considers a sum-over-states expression for the hyperpolarizabilities, we have in the denominator one has electronic state energies (which are identical for CC and EOMCC), while in the numerator appears the product of ground-to-excited state transition matrix elements (from linear response) and state to state transition matrix elements ( LABEL:SI-state-to-state-tm, given in terms of the contractions found in subsection 2.1). The consequence of the expression of the numerator is that, if the linear response terms for CC and EOMCC are similar while the full quadratic response function is not, the difference should be coming to the state to state transition matrix elements; or, in other words, in the description of transition moments between excited states.
@@
The dipole polarizabilities, shown in Figure 3 (values for dipole moments and polarizabilities are reported in LABEL:SI-tab:polar-hx-table) increase across the series, which is consistent with the increase in size of the halides accompanied by their decreasing electronegativity. The differences between CC and EOMCC do increase across the series for the and components and consequently, for the mean polarizability as well. However, in contrast to the hyperpolarizabilities, in relative terms these differences are roughly the same and of about 2% for and 1.5% for .
@@
The difference between CC and EOMCC is more strongly felt for the anisotropy, which is about 3% for HF, but 8-12% for HCl to HAt, and reaches about 30% for HTs, though in absolute terms the differences are still modest. An intriguing observation is that the polarizability anisotropy, while becoming less important for the almost spherical HAt and HTs molecules, changes sign upon going from HAt to HTs. This means that HTs can be more easily polarized in the direction perpendicular to the bond while the reverse is true for the lighter homologues that are more easily polarized along the bond. The sign change occurs with spin-orbit coupling but not without (though in magnitude the spin-orbit and spin-free anisotropies are very close to each other), making further studies necessary to better characterize it. We note that our CCSD dipole moments compare well with available experimental NISTconstants, Hohm2013 or theoretical work (for HTs) thierfelder_scalar_2009, though for reaching benchmark accuracy we would need to account for higher excitations, something which is also beyond the scope of this work. These findings add to the body of data supporting the argument that differences between LR-CC and LR-EOMCC properties are in practice small, and in particular when correlating the same number of electrons.
As mentioned above, another consequence of the similarity between CC and EOMCC polarizabilities is that it suggests their difference arises from the state to state transition matrix elements. While at this time we have not yet implemented the calculation of excited state properties (and such implementation is beyond the scope of this work), we can nevertheless gain some insight by considering the individual contributors to the value of the QR-CC functions.
@@
We see that across the halide series the dominant contributions typically arise from the terms from eq. subsection 2.1, involving both the perturbed amplitudes and multipliers ( and ). Furthermore, we observe that these two contributions to the final value in general tend to counteract each other, and that the magnitude of the “doubles” contributions increases as the halide becomes heavier. Terms from eq. subsection 2.1, involving only perturbed amplitudes and ground-state multipliers ( and ) show smaller (but non-negligible) net contributions than those above. They also increase in absolute value as the halide becomes heavier.
@@
The relative importance of contribution of the terms in eq. subsection 2.1 also changes depending on which response function we look at. For , the term is the one which contributes the most, but for there are more important cancellations between the “singles” and “doubles” contributions from it so that the term becomes a more important net contributor, being the dominant one for the HI molecule followed by , whereas and show similar contributions.
@@
4.2 Verdet constant of noble gases (Xe-Og) and hydrogen halides (HF-HI)